OpenStudy (anonymous):

R(t)=30t+525/165t+7000 where t is the nuimber of years since 1960....will t be=to 53?

jimthompson5910 (jim_thompson5910):

what is your question exactly? you want to know R(t) when t = 53?

OpenStudy (anonymous):

Use the following scenario to answer question 4 and 5. (HINT: a ratio can be written as a percentage) According to the Federal Buerau of Investigations the ratio of violent crimes to the total crimes committed in the U.S. is given by the following function: R(t)=30t+525/165t+7000 where t is the number of years since 1960...State the meaning of the horizontal asymptote as it relates to the stated scenario.

OpenStudy (anonymous):

a word problem... so i thought t=53 if the question addresses this year (2013) but I am not good with word problems, so i think the horizontal asymptote is 30/165...just dont know what is next

jimthompson5910 (jim_thompson5910):

ok so this is one big fraction I'm assuming

jimthompson5910 (jim_thompson5910):

if so, then as t gets larger and larger, then R(t) will slowly start to approach 30/165 = 2/11

jimthompson5910 (jim_thompson5910):

Basically this means that as time progresses into the future (much further into the future), the ratio of violent crimes will level out at 2/11

jimthompson5910 (jim_thompson5910):

I'm not sure why you are using t = 53 though

OpenStudy (anonymous):

i just thought 53 will be number of years since 1960- but i was not sure if this is what i had to solve

jimthompson5910 (jim_thompson5910):

oh no, you just think very far into the future (say 1000 years) and basically the ratio will approach or be close enough to 2/11 and this is the horizontal asymptote

OpenStudy (anonymous):

thank you so much! will the vertical intercept be 7000?

jimthompson5910 (jim_thompson5910):

no you plug in t = 0 and then evaluate

OpenStudy (anonymous):

ok ~ :)

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

0.7745

OpenStudy (anonymous):

also got 0.181818

jimthompson5910 (jim_thompson5910):

you should have gotten 525/7000 = 3/40 = 0.075 check your work again

OpenStudy (anonymous):

.075 you are right

OpenStudy (anonymous):

thank you for all your help!

jimthompson5910 (jim_thompson5910):

sure thing, I'm glad I was helpful